# Towards Graph Foundation Models: A Survey and Beyond

### v3

@@ 3.2.2 Intrinsic Differences
The intrinsic differences between GFM and LLM primarily manifest in two aspects: data and tasks. As for input data, language foundation models are primarily designed for processing Euclidean data, i.e., text. They are trained on vast text corpora, which are inherently sequential and follow a linear order of words or tokens. Graph foundation models, on the other hand, are designed to handle non-Euclidean data (represented as graph structures) or a mixture of Euclidean data (like graph attributes) and non-Euclidean data. Compared to text data, graph data can capture complex data relationships and is typically more sparse. Moreover, as mentioned in Section 3.1.3, different graphs may exhibit significant differences in type or scale, all of which pose challenges in the design of GFMs. Furthermore, language data, even when sourced from texts in different domains, still share a common vocabulary. On the other hand, different graph data may lack this common foundation. For instance, nodes represent atoms in a molecular graph, while nodes represent users in a social network graph, which are entirely different.
@@ 3.2.3 Extrinsic Differences
Backbone Architectures. Language foundation models, such as GPT-3 [50] and LLaMA [53], are mostly based on the Transformer architecture. The advantages of Transformer in terms of expressive power, scalability, parallelizability, and its excellent performance in handling various NLP tasks have made it the mainstream backbone architecture for LLMs. However, for GFMs, using mainstream GNNs as the backbone architecture may not necessarily be suitable. This is mainly because the expressive power and generalization of GNNs have limitations, and their parameter sizes are often too small to exhibit emergent abilities. Despite recent research efforts to design Graph Transformers [74] or models that incorporate LLMs [91], there is still no unified backbone architecture for GFMs.
Homogenization. Language foundation models are relatively easy to homogenize. This means that various NLP tasks can be formulated as the same task [92], making it possible to use a single model with unified training paradigm for a wide range of tasks. However, due to the poor transferability of graph structural knowledge, homogenization is more challenging for GFMs. Existing work attempts to achieve homogenization by unifying various tasks as link prediction [65] or graph-level tasks [66]. Additionally, constructing a data-task heterogeneous graph [70] may establish connections between different tasks, but it is a more complex process.
Domain Generalization. Language foundation models have demonstrated strong domain generalization capabilities. They can often perform well on tasks and datasets that were not seen during training, showcasing their ability to generalize across various language-related domains. However, due to the diversity and lack of unified vocabulary of graph data, GFMs generally exhibit weaker generalization across datasets, especially when moving to cross-domain graph data [86]. Their performance may degrade significantly when faced with graph structures or characteristics that differ substantially from their training data. Achieving strong domain generalization remains a challenging research problem for GFMs.
Emergence. Language foundation models have shown emergent abilities, where they can generate coherent and contextually relevant text based on few examples or instructions. Representative emergent abilities include in-context learning [67], chain of thought reasoning [68] and zero-shot generation [69]. However, GFMs have not demonstrated obvious emergent abilities to the same extent as language foundation models. Only a few recent studies discuss the in-context learning [70], graph reasoning [71] and zero-shot graph generation [72] abilities of GFMs.
@@ 3.3 Summary
In this section, we define the concept of graph foundation models and related technologies, and compares graph foundation models with language foundation models. In the following sections, we will introduce three categories of methods for implementing graph foundation models, along with representative works for each method, as illustrated in Figure 2. GNN-based models use GNN as the backbone architecture, while LLM-based models transform the graph into the input format of LLM and use LLM as the backbone architecture. GNN+LLM-based models, on the other hand, utilize both GNN and LLM as the backbone architecture simultaneously. The distinction in backbone architecture also impacts the methods for pre-training and adaptation. Therefore, in the following sections, we will introduce the backbone architectures, pre-training, and adaptation strategies for each category of methods, seperately.
@@ 4 GNN-based Models
{forest}
forked edges, folder indent=1cm, where=level()¡1folder, grow’=east, where=level()¿0l sep+=1cm, for tree= fork sep=4mm, thick, edge=thick, font=, if n children=0if n=1yshift=-5mm, for parent=s sep=0mmdraw, minimum height=4ex, minimum width=4cm [Towards Graph Foundation Models, calign=edge midpoint, s sep=2cm, font=[GNN-based Models [Backbone Architectures [Message Passing-based 4.1.1] [Graph Transformer-based 4.1.2] [, no edge, for children=no edge] ] [Pre-training [Contrastive Methods 4.2.1] [Generative Methods 4.2.2] [, no edge, for children=no edge] ] [Adaptation [Fine-Tuning 4.3.1] [Prompt-Tuning 4.3.2] ] ] [LLM-based Models [Backbone Architectures [Graph-to-Token 5.1.1] [Graph-to-Text 5.1.2] [, no edge, for children=no edge] ] [Pre-training [Language Modelling 5.2.1] [Masked Language Modelling 5.2.2] [, no edge, for children=no edge] ] [Adaptation [Manual Prompting 5.3.1] [Automatic Prompting 5.3.2] ] ] [GNN+LLM-based Models [Backbone Architectures [GNN-centric 6.1.1] [Symmetric 6.1.2] [LLM-centric 6.1.3] ] [Pre-training [GNN or LLM-based 6.2.1] [Alignment-based 6.2.2] [, no edge, for children=no edge] ] [Adaptation [Fine-Tuning 6.3.1] [Prompt-Tuning 6.3.2] ] ] ]
Figure 2: Taxonomy of existing works towards graph foundation models.
Thanks to effective model architectures and training paradigms, language models have achieved remarkable performance in natural language processing tasks. The backbone, pre-training and adaptation techniques employed in language models have inspired a series of corresponding efforts in the field of graph-based tasks. In this section, we will delve into GNN-based models, which draw inspiration from the model architectures or training paradigms used in NLP to apply them to graph tasks. Importantly, unlike the LLM-based models and GNN+LLM-based models to be introduced in the following sections, GNN-based models do not explicitly model text data in their pipeline. We have summarized and categorized the works mentioned in this section in Table II.

### v4

@@ 3.3.2 Intrinsic Differences
The intrinsic differences between GFM and LLM primarily manifest in two aspects: data and tasks. As for input data, language foundation models are primarily designed for processing Euclidean data, i.e., text. They are trained on vast text corpora, which are inherently sequential and follow a linear order of words or tokens. On the other hand, GFMs are designed to handle non-Euclidean data (represented as graph structures) or a mixture of Euclidean data (like graph attributes) and non-Euclidean data. Compared to text data, graph data can capture complex data relationships and is typically more sparse. Moreover, different graphs may exhibit significant differences in type or structure/geometry, all of which pose challenges in the design of GFMs. Furthermore, language data, even when sourced from texts in different domains, still share a common vocabulary. On the other hand, different graph data may lack this common foundation. For instance, nodes represent atoms in a molecular graph, while nodes represent users in a social network, which are entirely different. Furthermore, graphs from different domains can have different structures. Some graphs have a more hierarchical structure, while others may have a more cyclical structure [70]. Moreover, for a single graph, the different regions can exhibit different structures. For example, In recommender systems, the user-user subgraph and item-item subgraph generally exhibit very distinct structures [71].
@@ 3.3.3 Extrinsic Differences
Backbone Architectures. LLMs, such as GPT-3 [73] and LLaMA [74], are mostly based on the Transformer architecture. The advantages of Transformer in terms of expressive power, scalability, parallelizability, and its excellent performance in handling various NLP tasks have made it the mainstream backbone architecture for LLMs. However, for GFMs, using mainstream GNNs as the backbone architecture may not necessarily be suitable. This is mainly because the expressive power and generalization of GNNs have limitations, and their parameter sizes are often too small to exhibit emergent abilities. To enhance the expressiveness of capturing the graph structure, recent works [70, 75, 71] efforts to extend GNN to mixed curvature Riemannian space, or to design graph transformers [63] or models that incorporate LLMs [76]. However, there is still no unified backbone architecture for GFMs.
Homogenization. LLMs are relatively easy to homogenize. This means that various NLP tasks can be formulated as the same task [77], making it possible to use a single model with unified training paradigm for a wide range of tasks. However, due to the poor transferability of graph structural knowledge, homogenization is more challenging for GFMs. Existing works attempt to achieve homogenization by unifying various tasks as link prediction [55] or graph-level tasks [56]. Additionally, constructing a data-task heterogeneous graph [59] may establish connections between tasks, but it is a more complex process.
Domain Generalization. LLMs have demonstrated strong domain generalization capabilities. They can often perform well on tasks and datasets that were not seen during training, showcasing their ability to generalize across various language-related domains. However, due to the diversity and lack of unified vocabulary of graph data, GFMs generally exhibit weaker generalization across datasets, especially when moving to cross-domain graph data [78]. Their performance may degrade significantly when faced with graph structures or characteristics that differ substantially from their training data. Achieving strong domain generalization remains a challenging research problem for GFMs.
Emergence. LLMs have shown emergent abilities, where they can generate coherent and contextually relevant text based on few examples or instructions. Representative emergent abilities include in-context learning [57], chain of thought reasoning [58] and zero-shot generation [79]. However, GFMs have not demonstrated obvious emergent abilities to the same extent as language foundation models. Only a few recent studies discuss the in-context learning [59], graph reasoning [60] and zero-shot graph generation [61] abilities of GFMs.
@@ 3.4 Summary
In this section, we define the concept of graph foundation models and related technologies, and compares graph foundation models with language foundation models. If readers wish to have a more comprehensive understanding of the concept of GFMs, they can refer to our supplementary materials A and B, where we introduce the impact of graph data and graph tasks on GFMs. In the following sections, we will introduce three categories of methods for implementing graph foundation models, along with representative works for each method. Specifically, GNN-based models use GNN as the backbone architecture, while LLM-based models transform the graph into the input format of LLM and use LLM as the backbone architecture. GNN+LLM-based models, on the other hand, utilize both GNN and LLM as the backbone architecture simultaneously. The distinction in backbone architecture also impacts the methods for pre-training and adaptation. Therefore, in the following sections, we will introduce the backbone architectures, pre-training, and adaptation strategies for each category of methods, seperately.
@@ 4 GNN-based Models
Thanks to effective model architectures and training paradigms, language models have achieved remarkable performance in natural language processing tasks. The backbone, pre-training and adaptation techniques employed in language models have inspired a series of corresponding efforts in the field of graph-based tasks. In this section, we will delve into GNN-based models, which draw inspiration from the model architectures or training paradigms used in NLP to apply them to graph tasks. Importantly, unlike the LLM-based models and GNN+LLM-based models to be introduced in the following sections, GNN-based models do not explicitly model text data in their pipeline. We have summarized and categorized the works mentioned in this section in supplemental material C.